# Pre-DPO: Improving Data Utilization in Direct Preference Optimization Using a Guiding Reference Model

### v2

@@ Abstract
Direct Preference Optimization (DPO) simplifies reinforcement learning from human feedback (RLHF) for large language models (LLMs) by directly optimizing human preferences without an explicit reward model. We find that during DPO training, the reference model plays the role of a data weight adjuster. However, the common practice of initializing the policy and reference models identically in DPO can lead to inefficient data utilization and impose a performance ceiling. Meanwhile, the lack of a reference model in Simple Preference Optimization (SimPO) reduces training robustness and necessitates stricter conditions to prevent catastrophic forgetting. In this work, we propose Pre-DPO, a simple yet effective DPO-based training paradigm that enhances preference optimization performance by leveraging a guiding reference model. This reference model provides foresight into the optimal policy state achievable through the training preference data, serving as a guiding mechanism that adaptively assigns higher weights to samples more suitable for the model and lower weights to those less suitable. Extensive experiments on AlpacaEval 2.0 and Arena-Hard v0.1 benchmarks demonstrate that Pre-DPO consistently improves the performance of both DPO and SimPO, without relying on external models or additional data. [^1] Our code and data can be found at https://github.com/DtYXs/Pre-DPO.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has become a crucial paradigm in aligning large language models (LLMs) with human preferences and values (Christiano et al., 2017; Ouyang et al., 2022; Achiam et al., 2023). This approach involves training an explicit reward model to provide reward signals and performing reinforcement learning under a KL constraint by a reference model instantiated identically to the initial policy. Direct Preference Optimization (DPO) (Rafailov et al., 2023), as a representative of reference-based preference optimization methods (Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2024; Gorbatovski et al., 2024), transforms the objective of RLHF into direct preference training of the language model, enabling offline preference learning under the constraint of the reference model.
@@
Recently, it has been shown that a reference model is not necessary for achieving effective preference optimization. Simple Preference Optimization (SimPO) (Meng et al., 2024), as a representative of reference-free preference optimization methods (Xu et al., 2023; Xu et al., 2024; Hong et al., 2024; Meng et al., 2024), eliminates the need for a reference model and yields better performance and efficiency, though at the cost of an increased risk of catastrophic forgetting (Meng et al., 2024). Meanwhile, other studies have empirically demonstrated that DPO can benefit either from relaxing the constraints imposed by the reference model (Gorbatovski et al., 2024) or from stronger external reference models (Liu et al., 2024). However, decreasing reliance on a reference model imposes stricter practical requirements to ensure effective learning (Meng et al., 2024), and clear methodologies for obtaining an appropriate reference model are still lacking.
@@
Despite the empirical efforts, the role of the reference model in DPO and its impact on training dynamics remain insufficiently explored. We find that during DPO training, the reference model plays the role of a data weight adjuster (see Section 4.1). It adaptively tends to assign higher weights to data aligned with itself while reducing weights for conflicting data. However, due to the common practice in DPO of initializing the policy and reference models identically (Rafailov et al., 2023), as training progresses, the reference model increasingly constrains the policy model by penalizing deviations, potentially introducing a performance ceiling. Moreover, identical initialization of policy and reference models results in the nearly uniform weighting of training examples during the early training stages, which is in contrast to prior studies showing that assigning non-uniform weights to training data can lead to improved learning and performance (Lin et al., 2017; Ren et al., 2018; Shu et al., 2019).
@@
In light of the limitations of conventional reference models, we hypothesize that an effective reference model for DPO should (1) differ from the initial policy model and (2) provide insights into potential directions for policy improvement based on the preference data. We define this type of reference model as a guiding reference model, which can better support learning by transforming the role of the reference model from a static constraint into a guide with foresight (see Section 4.2). Building on this insight, we propose Pre-DPO, a simple yet effective training paradigm that enhances data utilization and improves the performance of existing preference optimization methods without relying on external models or additional data. Pre-DPO first optimizes the initial policy using a standard preference optimization method (e.g., DPO or SimPO). The resulting optimized policy is then employed as the guiding reference model. Finally, the initial policy is re-optimized using DPO with this guiding reference model, yielding a better-optimized policy through more effective data reweighting. The guiding reference model in Pre-DPO essentially serves as an adaptive guiding mechanism that dynamically assigns higher weights to samples more suitable for the model and lower weights to those less suitable. Specifically, these suitable cases typically correspond to examples that are easier to learn, allowing the model to leverage data that aligns well with its learning trajectory efficiently.
Figure 1: Pre-DPO further enhances the performance of DPO and SimPO by leveraging a guiding reference model, consistently across different models and scales.
@@
Our extensive experiments results conducted on the Llama3.2 (Grattafiori et al., 2024) and Qwen2.5 (Yang et al., 2024) model series demonstrate that Pre-DPO can consistently improve the performance of both DPO and SimPO on the AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v0.1 (Li et al., 2024) benchmarks, achieving average improvements of 2.5 points in the AlpacaEval 2 length-controlled win rate (LC) and 2.6 points in the Arena-Hard v0.1 win rate (WR). By introducing the guiding reference model, Pre-DPO can further improve the performance of existing well-tuned preference optimization methods (as illustrated in Figure 1), effectively overcoming the performance ceiling caused by inefficient data utilization under traditional reference model settings. Notably, Pre-DPO does not rely on external models or additional data, making it highly flexible and easy to deploy.
@@ 2 Related work
RLHF is currently an effective technique for aligning LLMs with human values and preferences (Ouyang et al., 2022; Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024; Shen et al., 2025). Typically, pretrained LLMs first undergo SFT to acquire the ability to follow human instructions. Subsequently, preference data is annotated by human or AI annotators to train a reward model. Finally, policy optimization is performed using reinforcement learning methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), Group Relative Policy Optimization (GRPO) (Shao et al., 2024), REINFORCE++ (Hu, 2025), etc. Although RLHF has demonstrated superior performance, the entire pipeline is complex and challenging.
@@
In contrast, offline preference optimization eliminates the need for an explicit reward model and avoids the complex online reinforcement learning optimization process. A well-optimized model obtained via preference optimization can also serve as a strong initial policy for subsequent reinforcement learning optimization (Yang et al., 2024). Depending on whether or not a reference model is required, preference optimization can be classified into two categories: reference-based preference optimization methods (Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2024; Gorbatovski et al., 2024) and reference-free preference optimization methods (Xu et al., 2023; Xu et al., 2024; Hong et al., 2024; Meng et al., 2024). DPO (Rafailov et al., 2023), as a representative of reference-based preference optimization methods, directly trains on preference data and implicitly optimizes the same objective as existing reinforcement learning algorithms. SimPO(Meng et al., 2024), as a representative of reference-free preference optimization methods, removes the need for a reference model and can achieve better results than DPO at the cost of lower training robustness (Meng et al., 2024). In this paper, we experimentally demonstrate that DPO can also benefit from reference-free preference optimization methods through guiding reference models.
@@
For the reference models in DPO, prior work (Liu et al., 2024) empirically suggested that DPO can benefit from a stronger and more suitable reference model in certain cases. However, it mainly focuses on stronger external reference models and does not provide a theoretical explanation for why they can be beneficial. Gorbatovski et al. (2024) proposed a dynamic update strategy to reset the reference model based on the current policy, which could weaken the regularization effect of the reference model and tends to assign more identical weights to the data samples with more frequent updates. In this work, we introduce the concept of a guiding reference model, analyze its impact on maximizing data utilization in DPO through data reweighting, and propose an effective methodology for leveraging a guiding reference model in DPO.
@@ 3 Preliminaries
Reinforcement learning from human feedback (RLHF).
@@
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\pi_{\theta}}}\left[r(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y\mid x)\parallel\pi_{\text{ref}}(y\mid x)\right],$ (1)
where $\mathcal{D}_{\pi_{\theta}}$ denotes the distribution of the policy model. The KL divergence term constrains the deviation of the policy $\pi_{\theta}$ from the reference model, and $\beta$ controls the strength of this constraint.
Direct Preference Optimization (DPO).
DPO (Rafailov et al., 2023) is a widely used reference-based preference optimization method that eliminates the need for explicit reward signals and has become a component in the post-training pipeline of many popular open-source LLMs (Bi et al., 2024; Jiang et al., 2024; Yang et al., 2024; Xu et al., 2025). It reformulates Eq. 1 into a direct optimization process on a preference dataset $\mathcal{D}=\{(x_{i},y^{+}_{i},y^{-}_{i})\}_{i=1}^{|\mathcal{D}|}$, where $x$ is the prompt, $y_{i}^{+}$ is the preferred response, and $y_{i}^{-}$ is the less-preferred response. The objective of DPO is as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\Big{)}\right],$ (2)
@@
A large $\beta$ imposes a strong constraint, which can limit the model’s performance improvement, while a small $\beta$ may result in insufficient constraints, leading to model degradation (Liu et al., 2024). In practical LLM training, $\pi_{\theta}$ and $\pi_{\text{ref}}$ are typically initialized as the supervised fine-tuning (SFT) model. $\pi_{\text{ref}}$ remains fixed during training.
Simple Preference Optimization (SimPO).
SimPO (Meng et al., 2024) is a reference-free preference optimization method that removes the reference model in DPO while introducing length normalization and target reward margin. Its loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y^{+}|}\log\pi_{\theta}(y^{+}\mid x)-\frac{\beta}{|y^{-}|}\log\pi_{\theta}(y^{-}\mid x)-\gamma\right)\right],$ (3)
where $|y^{+}|$ and $|y^{-}|$ denote the lengths, $\beta$ is a hyperparameter constant, and $\gamma$ is the target reward margin. SimPO has the potential to surpass DPO in performance but suffers from reduced robustness due to the lack of reference constraints (Meng et al., 2024).
@@ 4 Method
Figure 2: An overview of Pre-DPO. DPO constrains training using the initial policy model as the reference, and SimPO is reference-free. Pre-DPO first obtains an optimized policy model via DPO or SimPO, resets it as a guiding reference model, and then re-optimizes the initial policy using DPO. This process improves data utilization and leads to a better optimized policy model.
In this section, we first present the motivation behind Pre-DPO, focusing on the limitations of the reference setting in DPO and the specific challenges Pre-DPO addresses. Then, we provide a detailed explanation of Pre-DPO, outlining how it adaptively reweights the training examples and describing its overall process (as shown in Figure 2).
4.1 The reference model in DPO is a data weight adjuster
From the loss function of DPO (Eq. 2), we can derive the gradient with respect to the parameters $\theta$:
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\beta\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\lambda\cdot\nabla_{\theta}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\theta}(y^{-}\mid x)}\right],$ (4)
@@ 4.1 The reference model in DPO is a data weight adjuster
When the policy $\pi_{\theta}$ and the reference $\pi_{\text{ref}}$ are initialized from an identical SFT model, $\lambda$ starts around the constant 0.5 in the early training period due to $\sigma(0)=0.5$. However, a constant $\lambda$ can lead to degeneracy (Rafailov et al., 2023), and more importantly, previous research in the field of example reweighting has shown that assigning appropriate and varying weights to training samples can improve model performances and data efficiency (Lin et al., 2017; Ren et al., 2018; Shu et al., 2019).
On the other hand, as training progresses, the reference continuously constrains the policy’s deviation by adjusting the value of $\lambda$. Specifically, when $\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}$ is large, it encourages a larger value of $\lambda$, promoting learning from the corresponding preference pair. A small ratio typically results in a reduced value of $\lambda$, which can reduce the model’s learning from that sample. Therefore, a suboptimal configured reference model can lead to suboptimal weighting of training samples.

### v3

@@ Abstract
Direct Preference Optimization (DPO) simplifies reinforcement learning from human feedback (RLHF) for large language models (LLMs) by directly training on offline preference data to align with human preferences. During DPO training, the reference model serves as a data weight adjuster. However, the common practice of initializing the policy and reference models identically in DPO can lead to inefficient data utilization and impose a performance ceiling. Meanwhile, the absence of a reference model in Simple Preference Optimization (SimPO) reduces training robustness and requires stricter conditions to prevent catastrophic forgetting. In this work, we propose Pre-DPO, a simple yet effective DPO-based training paradigm that improves preference optimization by introducing a guiding reference model. This reference model provides foresight into the desired policy state achievable through the training preference data, serving as a guiding mechanism that adaptively assigns higher weights to samples more suitable for the model and lower weights to those less suitable. Extensive experiments on the AlpacaEval 2 and Arena-Hard v0.1 benchmarks demonstrate that Pre-DPO consistently improves the performance of both DPO and SimPO, without relying on external models or additional data.
Codehttps://github.com/DtYXs/Pre-DPO
@@ 1 Introduction
Preference-based training has become a widely adopted and effective paradigm for aligning large language models (LLMs) with human values and preferences. Direct Preference Optimization (DPO) (rafailov2023direct), as a representative of reference-based preference optimization methods (rafailov2023direct; ethayarajh2024kto; azar2024general; gorbatovski2025learn), directly trains LLMs on preference data under the constraint of a reference model, without relying on an explicit reward model or complex online reinforcement learning.
Figure 1: Pre-DPO introduces a guiding reference model derived from the optimized policy to guide re-optimization, transforming the reference from a constraint into an informed guide with foresight.
@@
Recently, it has been shown that a reference model is not necessary for achieving effective preference optimization. Simple Preference Optimization (SimPO) (meng2024simpo), as a representative of reference-free preference optimization methods (xu2023some; xu2024contrastive; hong2024orpo; meng2024simpo; nath2025dpl), eliminates the need for a reference model and yields better performance and efficiency, though at the cost of an increased risk of catastrophic forgetting (meng2024simpo). Meanwhile, other studies have empirically demonstrated that DPO can benefit either from relaxing the constraints imposed by the reference model (gorbatovski2025learn) or from stronger external reference models (liu2024understanding). However, decreasing reliance on a reference model imposes stricter practical requirements to ensure effective learning, and clear methodologies for obtaining an appropriate reference model are still lacking.
@@
Despite the empirical efforts, the role of the reference model in DPO and its impact on training dynamics remain insufficiently explored. During DPO training, the reference model serves as a data weight adjuster (see Section 4.1). It adaptively tends to assign higher weights to data aligned with itself while reducing weights for conflicting data. However, due to the common practice in DPO of initializing the policy and reference models identically (rafailov2023direct), as training progresses, the reference model increasingly constrains the policy model by penalizing deviations, potentially introducing a performance ceiling. Moreover, identical initialization of policy and reference models results in the nearly uniform weighting of training examples during the early training stages, which is in contrast to prior studies showing that assigning non-uniform weights to training data can lead to improved learning and performance (lin2017focal; ren2018learning; shu2019meta).
@@
In light of the limitations of conventional reference models, we hypothesize that an ideal reference model for DPO should originate from the initial policy model and provide insights into potential directions for policy improvement based on the preference data. We define this type of reference model as a guiding reference model, which can better support learning by transforming the role of the reference model from a constraint into a guide with foresight (see Section 4.2). Building on this insight, we propose Pre-DPO, a simple yet effective training paradigm that enhances data utilization and improves the performance of existing preference optimization methods without relying on external models or additional data, as shown in Figure 1. Pre-DPO first optimizes the initial policy using a standard preference optimization method (e.g., DPO or SimPO). The resulting optimized policy is then employed as the guiding reference model. Finally, the initial policy is re-optimized using DPO with this guiding reference model, yielding a better-optimized policy through more effective data reweighting. The guiding reference model in Pre-DPO essentially serves as an adaptive guiding mechanism that naturally assigns higher weights to samples more suitable for the model and lower weights to those less suitable. In practice, these suitable cases typically correspond to examples that are easier to learn, allowing the model to efficiently leverage data that aligns well with its learning trajectory.
@@
We evaluate Pre-DPO on the Llama3.2-3B (grattafiori2024llama) and Qwen2.5-7B (yang2024qwen2) model series across AlpacaEval 2 (alpaca_eval; dubois2024length) and Arena-Hard v0.1 (li2024crowdsourced) benchmarks. The experimental results show that Pre-DPO consistently improves the performance of both DPO and SimPO, achieving average gains of 2.5 points in length-controlled win rate (LC) on AlpacaEval 2 and 2.6 points in win rate (WR) on Arena-Hard v0.1. By introducing the guiding reference model, Pre-DPO can further improve the performance of existing well-tuned preference optimization methods, effectively overcoming the performance ceiling caused by inefficient data utilization under traditional reference model settings. Notably, Pre-DPO does not rely on external models or additional data, making it highly flexible and easy to deploy.
@@ 2 Related Work
Reinforcement learning from human feedback (RLHF) (christiano2017deep; stiennon2020learning) has become an effective approach for aligning LLMs with human values and preferences (ouyang2022training; achiam2023gpt; grattafiori2024llama; yang2024qwen2; shen2025exploring). Typically, pretrained LLMs first undergo supervised fine-tuning (SFT) to learn instruction-following behavior. Subsequently, reinforcement learning is conducted using external reward models and policy optimization algorithms (schulman2017proximal; shao2024deepseekmath; hu2025reinforce++), which are typically performed online. While RLHF has demonstrated strong performance, its full pipeline remains complex and resource-intensive. DPO (rafailov2023direct) simplifies the RLHF process by directly optimizing on preference data in an offline setting. In this work, we focus on improving data utilization in such offline preference optimization.
@@
Offline preference optimization eliminates the need for an explicit reward model and avoids the complex online learning optimization process. A well-optimized model obtained via offline preference optimization can also serve as a strong initial policy for subsequent online reinforcement learning optimization (yang2024qwen2). Depending on whether or not a reference model is required, offline preference optimization can be classified into two categories: reference-based preference optimization methods (rafailov2023direct; ethayarajh2024kto; azar2024general; kim2025sdpo; gorbatovski2025learn) and reference-free preference optimization methods (xu2023some; xu2024contrastive; hong2024orpo; meng2024simpo; nath2025dpl). DPO (rafailov2023direct), as a representative of reference-based preference optimization methods, directly trains on preference data and implicitly optimizes the same objective as existing reinforcement learning algorithms. SimPO (meng2024simpo), as a representative of reference-free preference optimization methods, removes the need for a reference model and can achieve better results than DPO at the cost of lower training robustness (meng2024simpo). In this paper, we experimentally demonstrate that DPO can also benefit from reference-free preference optimization methods by leveraging a guiding reference model.
@@
For the reference models in DPO, prior work (liu2024understanding) empirically suggested that DPO can benefit from a stronger and more suitable reference model in certain cases. However, it mainly focuses on stronger external reference models and does not provide a theoretical explanation for why they can be beneficial. gorbatovski2025learn proposed a dynamic update strategy to reset the reference model based on the current policy, which could weaken the regularization effect of the reference model and tends to assign more identical weights to the data samples with more frequent updates. In this work, we introduce the concept of a guiding reference model, analyze its role in enhancing data utilization in DPO through better data reweighting, and propose an effective methodology for leveraging it in the DPO framework.
@@ 3 Preliminaries
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\mathcal{D}\times\pi_{\theta}}\left[r(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y|x)\parallel\pi_{\text{ref}}(y|x)\right],$ (1)
where $\mathcal{D}\times\pi_{\theta}$ denotes the joint distribution of the prompt $x$ and the response $y$ from $\pi_{\theta}(y\mid x)$. The KL divergence term constrains the deviation of the policy $\pi_{\theta}$ from the reference model, and $\beta$ controls the strength of this constraint.
DPO (rafailov2023direct) is a widely used reference-based preference optimization method that eliminates the need for explicit reward signals and has become a component in the post-training pipeline of many popular open-source LLMs (bi2024deepseek; jiang2024mixtral; yang2024qwen2; xu2025qwen25omni). It reformulates Eq. 1 into a direct optimization process on a preference dataset $\mathcal{D}=\{(x_{i},y^{+}_{i},y^{-}_{i})\}_{i=1}^{|\mathcal{D}|}$, where $x$ is the prompt, $y_{i}^{+}$ is the preferred response, and $y_{i}^{-}$ is the less-preferred response. The objective of DPO is as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})\!=\!$ $-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\Big(\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}$
$\qquad\qquad-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\Big)\Big]$ (2)
@@
A large $\beta$ imposes a strong constraint, which can limit the model’s performance improvement, while a small $\beta$ may result in insufficient constraints, leading to model degradation (liu2024understanding). In practical LLM training, $\pi_{\theta}$ and $\pi_{\text{ref}}$ are typically initialized as the supervised fine-tuning (SFT) model. $\pi_{\text{ref}}$ remains fixed during training.
SimPO (meng2024simpo) is a reference-free preference optimization method that removes the reference model in DPO while introducing length normalization and target reward margin. Its loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=$ $-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\Big[\log\sigma\Big($
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\beta}{|y^{+}|}\log\pi_{\theta}(y^{+}\mid x)-\frac{\beta}{|y^{-}|}\log\pi_{\theta}(y^{-}\mid x)-\gamma\Big)\Big]$ (3)
where $|y^{+}|$ and $|y^{-}|$ denote the lengths, $\beta$ is a hyperparameter constant, and $\gamma$ is the target reward margin. SimPO has the potential to surpass DPO in performance but suffers from reduced robustness due to the lack of reference constraints (meng2024simpo).
@@ 4 Method
Figure 2: An overview of Pre-DPO. DPO constrains training using the initial policy model as the reference, while SimPO is reference-free. Pre-DPO first optimizes a policy model using DPO or SimPO, then resets it as a guiding reference model, and re-optimizes the initial policy using DPO. This process enhances data utilization and results in a better-optimized policy model.
In this section, we first present the motivation behind Pre-DPO, focusing on the limitations of the reference setting in DPO and the specific challenges Pre-DPO addresses. Then, we provide a detailed explanation of Pre-DPO, outlining how it adaptively reweights the training examples and describing its overall process.
4.1 Challenges in Data Reweighting with Vanilla DPO’s Reference Model
From the loss function of DPO (Eq. 2), we can derive the gradient with respect to the parameters $\theta$ (the detailed derivation is provided in Appendix A):
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\beta\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\lambda\cdot\nabla_{\theta}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\theta}(y^{-}\mid x)}\right],$ (4)
@@ 4.1 Challenges in Data Reweighting with Vanilla DPO’s Reference Model
When the policy $\pi_{\theta}$ and the reference $\pi_{\text{ref}}$ are initialized from an identical SFT model, $\lambda$ starts around the constant 0.5 in the early stage of training due to $\sigma(0)=0.5$. However, a constant $\lambda$ can lead to degeneracy (rafailov2023direct), and more importantly, previous research in the field of example reweighting has shown that assigning appropriate and varying weights to training samples can improve model performances and data efficiency (lin2017focal; ren2018learning; shu2019meta).
On the other hand, as training progresses, the reference continuously constrains the policy deviation by adjusting the value of $\lambda$. Specifically, when $\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}$ is large, it encourages a larger value of $\lambda$, promoting learning from the corresponding preference pair. Conversely, a smaller ratio typically leads to a lower $\lambda$, which in turn reduces the influence of that sample on the learning process. The difference in $\lambda$ between DPO and Pre-DPO is largely driven by the reference model, particularly in the early stage of training. Therefore, a suboptimally configured reference model can result in suboptimal weighting of training samples.